# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ A.5.1.2 Random graphs
    Barabási-Albert graphs[barabasi1999emergence]: The Barabási-Albert graph is a class of random scale-free networks, which grows in size via the following mechanism. A new vertex is adjacent to $k$ old vertices such that the probability of choosing an old vertex depends on the degree of the old vertex. Therefore, the vertices with high degree are preferred while growing the graph, leading to fewer vertices with large vertex degree. In our analysis, we set $k=3$ and consider all graphs in this class with the number of vertices $5\leq N\leq 5009$. Two of the graphs from this family are presented in Fig. A.8(a).
@@
    Newman-Watts-Strogatz graphs [newman1999renormalization]: We first arrange $N$ vertices over a ring lattice. Each vertex in the ring is connected to its $g$ nearest neighbors or (g1)(g-1) nearest neighbors if $g$ is odd. Further we add new edges as follows. For each edge $(v_{i},v_{j})$ in the underlying “$N$-ring with $g$ nearest neighbors", with probability $p$ add a new edge $(v_{i},v_{k})$ with randomly-chosen existing node $v_{k}$. For our analysis, we set $g=3$ and $p=1$ with seed value $19$. We generate all Newman-Watts-Strogatz graphs with the number of vertices in the range $5\leq N\leq 5005$. Fig. A.8(d) contains two instances of this family.
@@
    Random regular graphs: We generate the random regular graphs using the algorithms discussed in [steger1999generating, kim2003generating]. Here, the product of the number of vertices $N$ and the regularity $k$ must be an even number. For our calculations, we fix $k=4$ and generate all graphs with the number of vertices $5\leq N\leq 5013$. Two graphs from the family are given in Fig. A.8(j).
@@
    Gn,p random graphs: We generate all Gn,p random graphs, in other words the Erdös-Rényi random graphs [erdds1959random], with the number of vertices $5\leq N\leq 5009$. The only parameter in this construction is $n$ which sets the number of vertices in the graph, i.e., $N=n$. Here, we set the probability of existence of an edge between two randomly selected vertices to be $p=0.8$. Two graphs of this family are given in Fig. A.9(bo).
@@
    Gaussian random partition graphs [brandes2003experiments]: A Gaussian random partition graph is constructed by generating $k$ partitions on the set of vertices $N$. The size of the partitions are drawn from a normal distribution with a fixed mean and variance. For our work, we consider the Gaussian random partition graphs with $k=5$ partitions whose size is determined by the Gaussian distribution with mean $5$ and variance $1$. The probability of establishing the edges between two vertices inside a partition is set to be $0.5$ and in between the partitions is fixed to be $0.4$. For our investigation, we generate all the Gaussian random partition graphs with vertices $5\leq N\leq 5009$. Two of the graphs from this family are given in Fig. A.9(p).
@@
    Geographical threshold graphs [masuda2005geographical, bradonjic2007giant]: To construct a graph from this family, we place all the $N$ vertices on a Cartesian plane, with the distances between the vertices chosen at random. Each vertex is assigned a random weight drawn from an exponential distribution with rate parameter $k=1$, in our case. Two vertices $p$ and $q$ with weights $x_{p}$ and $x_{q}$ are connected if $(x_{p}+x_{q})r^{-2}\geq 10$, where $r$ is the Euclidean distance between two vertices. We generate graphs of this family in the range $5\leq N\leq 5009$. Two of the graphs from this family are presented in Fig. A.9(s).
@@
    Soft random geometric graphs [penrose2016connectivity]: To construct the soft random geometric graph, we place $N$ vertices randomly on a Cartesian plane. A probability function determines the connection between two vertices, which takes the Euclidean distance between two vertices as an input. If the Euclidean distance, $r$, between two vertices is at most $1$, then the probability of joining them by an edge is decided by the probability density function $e^{-r}$. Here, we generate graphs with number of vertices $5\leq N\leq 5009$. Fig. A.9(v) presents two graphs from this family.
@@
    Thresholded random geometric graphs: To construct this graph family, we place $N$ vertices randomly on the Cartesian plane, and each vertex is assigned a weight that is randomly chosen from an exponential distribution of rate parameter $k$. We set $k=1$. Two vertices are connected by an edge if the sum of the their weights is greater than or equal to $2$ and the Euclidean distance between the vertices is less than or equal to $1$. We generate graphs with vertices $5\leq N\leq 5009$. Fig. A.9(y) displays two instances of this graph family.
@@
    Planted partition graphs [condon2001algorithms]: A planted partition graph has $N=l\times n$ vertices distributed into $l$ groups with $n$ vertices in each. Two vertices in the same group are linked with a probability $p$. Two vertices belonging to two different groups are linked with probability $q$. For our calculations, we consider $l=2$, $p=0.5$ and $q=0.4$. To generate a graph family, we vary $n$, such that, the number of vertices $N$ varies from $10$ to $5014$. Fig. A.9(ab) presents two instances from this graph family.
@@
    Random geometric graphs [penrose2003random]: We place $N$ vertices randomly on a Cartesian plane join two vertices by an edge if their Euclidean distance, $r$ is at most $1$. For our calculations, we construct a graph family where the the number of vertices range from $7$ to $5008$. Fig. A.9(ae) contains two instances of this graph family.
@@
    Uniform random intersection graphs [fill2000random]: These graphs are generated from random bipartite graphs. First, we construct a bipartite graph with two partite sets containing $N$ and N3N-3 vertices, respectively. The probability of existence of an edge between the partite sets is set to be $0.6$. Then, the bipartite graph is projected onto the partite sets having $N$ vertices. There is an edge between two vertices in the resultant graph, if those two vertices share a common neighbor in the first bipartite graph. We vary the number of vertices $N$ from $5$ to $2999$, due to computational limitations. Fig. A.9(ah) represents two graphs belonging to this family.
@@
    Random lobster graphs: This graph is a tree which generates a caterpillar graph when all the leaf vertices are removed. We vary the number of vertices present in the backbone of the graph to generate a family. An edge between two vertices belonging to the backbone of the graph is drawn with probability $0.6$, while the probability of edge creation beyond backbone level is set to be $0.5$. We set the seed value to $19$. The number of vertices in the graph varies in the range $20\leq N\leq 5014$. Fig. A.9(bl) contains two instances of random lobster graph family.
A.5.2  Graphs whose incidence matrices are considered in our studies
@@
A.5.2.1 Non-random directed graphs
@@
    Paley graphs: For an odd prime number $N$, the vertex set of the Paley graph is $\mathbb{Z}/N\mathbb{Z}$. A directed edge exists from vertex $u$ to $w$ if $u-w\equiv{x^{2}(\text{mod}\ N)}$, where $x^{2}\in\mathbb{Z}/N\mathbb{Z}$ and $N$ is the number of vertices. Also, to avoid bi-directed edges, $N\equiv 3(\text{mod}\ 4)$, so that $1\not\equiv x^{2}(\text{mod}\ N)$. We follow the construction given in [networkx]. Our system size $N^{\prime}$ varies from $6$ to $9730$. Fig. A.11(y) presents two instances of this graph family.
@@
    Directed hypercube graphs: Similar to the undirected hypercube graphs of dimension $n$, we consider the vertex set $V(G_{2}^{n})=\{0,1\}^{\times n}$ which contains $N=2^{n}$ vertices. There is a directed edge from vertex $v_{1}$ to $v_{2}$ if the Hamming distance between the vertices is $1$ and the Hamming weight of $v_{1}<v_{2}$. Here, Hamming weight of a vertex refers to the number of $1$s in the corresponding $n$-tuple. This arrangement of directed edges indicates that there is only one source and sink in a directed hypercube graph. The number of edges in a hypercube graph grows as $2^{n-1}n$. Hence, the system size in a directed hypercube graph grows as $N^{\prime}=2^{n}(1+n/2)\sim\mathcal{O}(2^{n}n)$. The number of vertices in the graph assists the graph-family to grow in terms of $2^{n}$, where $n=2,3,\cdots$. Our system size varies $N^{\prime}$ from $8$ to $131072$. Fig. 4(a) presents two graphs belonging to this family.
A.5.2.2 Random directed graphs
We consider nine random directed graphs in this article. For the sake of reproducibility, we fix seed value of every random directed graph family to be $19$. Further, we followed the constructions from Ref. [networkx] with slight modifications to avoid bi-directed edges and self-loops in the graphs.
@@ A.5.2.2 Random directed graphs
    GN graph: A GN graph grows the network by linking a newly added vertex to one of the existing vertices based on their degree. We define a linear function $f(d)=d$, where $d$ is the degree of the vertex. This function determines the probability of connecting the new vertex to one of the existing vertices. In our work, the value of system size $N^{\prime}$ goes from $9$ to $9997$ for GN graphs with source and sink vertices. Also, the system size varies from $11$ to $9998$ for GN graphs without source and sink vertices. Figs. A.13(a) and A.13(d) present two instances each for without and with source and sink vertices, respectively.
@@
    GNC graph [krapivsky2005network]: A GNC graph is grown by adding a directed edge from a new vertex to an existing vertex chosen randomly. Further, directed edges go from the new vertex to the successors of the existing vertex to which it is connected. A vertex, $v_{u}$, is said to be the successor to a vertex $v_{w}$, if there is a directed edge going from $v_{w}$ to $v_{u}$. For GNC graphs with source and sink vertices, we consider $N^{\prime}$ in the range of $11$ to $8498$. Also, for the family of GNC graphs without any source and sink vertices, the system size $N^{\prime}$ varies from $11$ to $6759$. Two GNC graphs without and with source and sink vertices are depicted in Figs. A.13(g) and A.13(j) respectively.
@@
    GNR graph [krapivsky2001organization]: The construction of the GNR graphs is similar to the GN graphs. Here, when a new vertex is added, a directed edge appears from the new vertex to a randomly chosen existing vertex, called the target vertex. There is a finite chance for this new directed edge to get redirected to first successor of target vertex. For our work, we set the probability of reconnection to $0.5$. For the GNR graphs with source and sink vertices, the system size grows from $9$ to $9997$. Also, for GNR graphs without source and sink vertices, $N^{\prime}$ varies from $12$ to $10000$. Figs. A.13(m) and A.13(p) respectively present the figures for graphs without and with source and sink vertices belonging to this family.
@@
    Gaussian random partition graphs: The underlying principles for the directed Gaussian random partition graph is similar to its undirected version as described in sub-section A.5.1.2. We follow the construction from Ref. [networkx] to build directed Gaussian random partition graphs. The values of the parameters to build this graph family are same as the values set in the undirected version. We vary $N^{\prime}$ from $12$ to $10897$ for both the cases. Figs. 5(a) and 5(d) have illustrations of two graphs each for without and with source and sink vertices of this family.
@@
    Planted partition graphs: Our construction for the planted partition graph is also similar to their undirected mentioned in the sub-section A.5.1.2. Here, we fix the number of vertices present in each partition to be $5$ and vary the number of partitions to grow the graph. We set the probability of linking vertices inside a partition to $0.8$ and between the partitions to $0.4$. For graphs with sink and source vertices, we vary $N^{\prime}$ from $228$ to $24405$. For graphs without sink and source vertices, we vary $N^{\prime}$ from $228$ to $20208$. Two graphs each for without and with source and sink vertices are presented in Figs. 5(g) and A.10(a) respectively.
@@
    Navigable small world graphs [kleinberg2000small]: The graph is defined on a two-dimensional $n\times n$ grid with $N=n^{2}$ vertices. The distance between two vertices $(a,b)$ and $(c,d)$ is defined as $\mathcal{R}=|a-c|+|b-d|$. Based on the distance between vertices, we define short-range and long-range connections. A vertex $(a,b)$ is connected to all vertices which are at distance $1$ to it. The number of long-range connections of a vertex is set to $1$. The probability of connecting the vertex to a target vertex at a distance $\mathcal{R}$ is proportional to $1/\mathcal{R}^{2}$. For graphs without source and sink vertices, $N^{\prime}$ is varied from $84$ to $20537$, whereas for graphs with source and sink vertices, $N^{\prime}$ goes from $83$ to $14827$. Figs. A.10(d) and A.10(g) show two graphs each for without and with source and sink vertices corresponding to this family.
@@
    Gn,p random graphs: Our construction for the Gn,p random graph is also similar to their counterparts mentioned in the sub-section A.5.1.2 with same values for the parameters. For graphs without sink and source vertices, $N^{\prime}$ is varied from $44$ to $42534$, while for graphs with sink and source vertices, $N^{\prime}$ ranges from $44$ to $21721$. Two graphs each for without and with source and sink vertices can be found in Figs. A.11(a) and A.11(d) respectively.
@@
    Random uniform $k$-out graphs: In this graph, $k$ directed edges go from every vertex of the graph to $k$ vertices chosen randomly without replacement. For our calculations, we set $k=2$. For graphs with source and sink, $N^{\prime}$ takes the values from $12$ to $10943$ and graphs without source and sink, $N^{\prime}$ goes from $13$ to $10817$. Figs. A.12(a) and A.12(d) present two graphs each for without and with source and sink vertices respectively.
@@
    Scale-free graphs [bollobas2003directed]: In this graph family, the in-degree and out-degree distributions of vertices follow power laws. We start with a cycle graph with 3 edges. A directed edge from a new vertex to an existing vertex, chosen according to its in-degree distribution, is added with probability $0.41$. An edge between two existing vertices are added with probability $0.54$ and the vertices are chosen according to their in-degree and out-degree distributions. The probability of adding a directed edge from an existing vertex, chosen according to its out-degree distribution, to a new vertex is $0.05$. The probability of choosing an existing vertex would depend on the in-degree (or out-degree) in addition to a constant bias. The constant bias for in-degree distribution is set to be $\delta_{\mathrm{in}}=0.2$ and for out-degree distribution is set to be $\delta_{\mathrm{out}}=0$. Here, we vary the system size $\mathcal{N}$ from $12$ to $10997$ for graphs without sink and source vertices. For graphs with sink and source vertices, $\mathcal{N}$ is varied from $11$ to $9995$. Figs. A.12(g) and A.12(j) present two instances each for graph without and with source and sink vertices respectively.
Appendix A.6 Beyond graph theoretic considerations: Other requirements and challenges
A.6.1  Algorithmic challenges
@@
A.6.2  Quantum hardware challenges
@@
    NISQ era requirements: Although HHL and other QLSs discussed in this work are outside of the scope of NISQ computers and would ideally require a fault-tolerant implementation, it is important to test these solvers on smaller system sizes for various applications even in the NISQ era, continuously pushing the boundaries as quantum hardware advances. A typical theme in carrying out such computations would be to rely upon classical resources (which do not necessarily scale well) to reduce quantum circuit depth and number of qubits, we term these gamut of techniques as resource reduction. In Section VII, we illustrated these ideas through toy graph instances. Our results show that with current quantum hardware capabilities, we can only probe $(4\times 4)$ size Laplacian matrices (unless we find special graphs and leverage techniques like all-qubit fixing, as explained in Section VII.0.2), indicating the actual gap between what is possible and what is our goal for the long-term. For perspective, it is routine to solve systems of linear equations involving matrices that are about $(10^{9}\times 10^{9})$ in size, for example, in quantum chemical computations on traditional computers [Timo].
@@
    Fault-tolerant quantum computing era requirements: Of most relevance in the late NISQ/ early fault-tolerant quantum computing eras would be graph families that provide exponential speed-up and polynomial speed-up (with a larger-than-quadratic degree polynomial), in view of overheads associated with quantum error corrected implementations eating up a sub-linear or a lower degree polynomial speed-up offered by good graph families and some better graph families (for example, see Ref. [prx2021], where the authors show that quadratic speedups will not enable quantum advantage on early generation fault-tolerant computers that use surface codes). It is hard to predict the extent to which such issues will plague good and better graph families in the fault-tolerant quantum computing era, where one can use codes with lower thresholds. However, in all of these cases, one aims at reducing either the runtime or the number of physical qubits or in an ideal case, both of them, in fault-tolerant QLS implementations. This would in turn depend on achieving lower physical error rates and faster gate times in the quantum hardware that is used, as well as choice of suitable quantum error correction codes. Each of these is formidable, and until these obstacles are cleared, truly realizing an advantage even with better and best graphs is hard.
Appendix A.7 Some theorems relevant to all-qubit fixing
Theorem A.7.1.
@@ Appendix A.7 Some theorems relevant to all-qubit fixing
$L\begin{bmatrix}0\\ 0\\ \cdots\\ 0\\ b_{i}\\ 0\\ \cdots\\ 0\\ b_{j}\\ 0\\ \cdots\\ 0\end{bmatrix}=\begin{bmatrix}l_{1i}-l_{1j}\\ l_{2i}-l_{2j}\\ \cdots\\ l_{i-1\ i}-l_{i-1\ j}\\ l_{ii}-l_{ij}\\ l_{i+1\ i}-l_{i+1\ j}\\ \cdots\\ l_{j-1\ i}-l_{j-1\ j}\\ l_{ji}-l_{jj}\\ l_{j+1\ i}-l_{j+1\ j}\\ \cdots\\ l_{Ni}-l_{Nj}\end{bmatrix}$. (A.13)

### v4

@@ Acknowledgements.
    Barabási-Albert graphs[6]: The Barabási-Albert graph is a class of random scale-free networks, which grows in size via the following mechanism. A new vertex is adjacent to $k$ old vertices such that the probability of choosing an old vertex depends on the degree of the old vertex. Therefore, the vertices with high degree are preferred while growing the graph, leading to fewer vertices with large vertex degree. In our analysis, we set $k=3$ and consider all graphs in this class with the number of vertices $5\leq N\leq 5009$. Two of the graphs from this family are presented in Fig. S.8(e).
@@
    Newman-Watts-Strogatz graphs [49]: We first arrange $N$ vertices over a ring lattice. Each vertex in the ring is connected to its $g$ nearest neighbors or (g1)(g-1) nearest neighbors if $g$ is odd. Further we add new edges as follows. For each edge $(v_{i},v_{j})$ in the underlying “$N$-ring with $g$ nearest neighbors", with probability $p$ add a new edge $(v_{i},v_{k})$ with randomly-chosen existing node $v_{k}$. For our analysis, we set $g=3$ and $p=1$ with seed value $19$. We generate all Newman-Watts-Strogatz graphs with the number of vertices in the range $5\leq N\leq 5005$. Fig. S.8(c) contains two instances of this family.
@@
    Random regular graphs: We generate the random regular graphs using the algorithms discussed in [61, 33]. Here, the product of the number of vertices $N$ and the regularity $k$ must be an even number. For our calculations, we fix $k=4$ and generate all graphs with the number of vertices $5\leq N\leq 5013$. Two graphs from the family are given in Fig. S.8(i).
@@
    Gn,p random graphs: We generate all Gn,p random graphs, in other words the Erdös-Rényi random graphs [20], with the number of vertices $5\leq N\leq 5009$. The only parameter in this construction is $n$ which sets the number of vertices in the graph, i.e., $N=n$. Here, we set the probability of existence of an edge between two randomly selected vertices to be $p=0.8$. Two graphs of this family are given in Fig. S.9(ao).
@@
    Gaussian random partition graphs [11]: A Gaussian random partition graph is constructed by generating $k$ partitions on the set of vertices $N$. The size of the partitions are drawn from a normal distribution with a fixed mean and variance. For our work, we consider the Gaussian random partition graphs with $k=5$ partitions whose size is determined by the Gaussian distribution with mean $5$ and variance $1$. The probability of establishing the edges between two vertices inside a partition is set to be $0.5$ and in between the partitions is fixed to be $0.4$. For our investigation, we generate all the Gaussian random partition graphs with vertices $5\leq N\leq 5009$. Two of the graphs from this family are given in Fig. S.9(g).
@@
    Geographical threshold graphs [45, 10]: To construct a graph from this family, we place all the $N$ vertices on a Cartesian plane, with the distances between the vertices chosen at random. Each vertex is assigned a random weight drawn from an exponential distribution with rate parameter $k=1$, in our case. Two vertices $p$ and $q$ with weights $x_{p}$ and $x_{q}$ are connected if $(x_{p}+x_{q})r^{-2}\geq 10$, where $r$ is the Euclidean distance between two vertices. We generate graphs of this family in the range $5\leq N\leq 5009$. Two of the graphs from this family are presented in Fig. S.9(i).
@@
    Soft random geometric graphs [51]: To construct the soft random geometric graph, we place $N$ vertices randomly on a Cartesian plane. A probability function determines the connection between two vertices, which takes the Euclidean distance between two vertices as an input. If the Euclidean distance, $r$, between two vertices is at most $1$, then the probability of joining them by an edge is decided by the probability density function $e^{-r}$. Here, we generate graphs with number of vertices $5\leq N\leq 5009$. Fig. S.9(k) presents two graphs from this family.
@@
    Thresholded random geometric graphs: To construct this graph family, we place $N$ vertices randomly on the Cartesian plane, and each vertex is assigned a weight that is randomly chosen from an exponential distribution of rate parameter $k$. We set $k=1$. Two vertices are connected by an edge if the sum of the their weights is greater than or equal to $2$ and the Euclidean distance between the vertices is less than or equal to $1$. We generate graphs with vertices $5\leq N\leq 5009$. Fig. S.9(m) displays two instances of this graph family.
@@
    Planted partition graphs [16]: A planted partition graph has $N=l\times n$ vertices distributed into $l$ groups with $n$ vertices in each. Two vertices in the same group are linked with a probability $p$. Two vertices belonging to two different groups are linked with probability $q$. For our calculations, we consider $l=2$, $p=0.5$ and $q=0.4$. To generate a graph family, we vary $n$, such that, the number of vertices $N$ varies from $10$ to $5014$. Fig. S.9(o) presents two instances from this graph family.
@@
    Random geometric graphs [52]: We place $N$ vertices randomly on a Cartesian plane join two vertices by an edge if their Euclidean distance, $r$ is at most $1$. For our calculations, we construct a graph family where the the number of vertices range from $7$ to $5008$. Fig. S.9(q) contains two instances of this graph family.
@@
    Uniform random intersection graphs [21]: These graphs are generated from random bipartite graphs. First, we construct a bipartite graph with two partite sets containing $N$ and N3N-3 vertices, respectively. The probability of existence of an edge between the partite sets is set to be $0.6$. Then, the bipartite graph is projected onto the partite sets having $N$ vertices. There is an edge between two vertices in the resultant graph, if those two vertices share a common neighbor in the first bipartite graph. We vary the number of vertices $N$ from $5$ to $2999$, due to computational limitations. Fig. S.9(s) represents two graphs belonging to this family.
@@
    Random lobster graphs: This graph is a tree which generates a caterpillar graph when all the leaf vertices are removed. We vary the number of vertices present in the backbone of the graph to generate a family. An edge between two vertices belonging to the backbone of the graph is drawn with probability $0.6$, while the probability of edge creation beyond backbone level is set to be $0.5$. We set the seed value to $19$. The number of vertices in the graph varies in the range $20\leq N\leq 5014$. Fig. S.9(am) contains two instances of random lobster graph family.
S.5.2  Graphs whose incidence matrices are considered in our studies
@@
S.5.2.1 Non-random directed graphs
@@
    Paley graphs: For an odd prime number $N$, the vertex set of the Paley graph is $\mathbb{Z}/N\mathbb{Z}$. A directed edge exists from vertex $u$ to $w$ if $u-w\equiv{x^{2}(\text{mod}\ N)}$, where $x^{2}\in\mathbb{Z}/N\mathbb{Z}$ and $N$ is the number of vertices. Also, to avoid bi-directed edges, $N\equiv 3(\text{mod}\ 4)$, so that $1\not\equiv x^{2}(\text{mod}\ N)$. We follow the construction given in [28]. Our system size $N^{\prime}$ varies from $6$ to $9730$. Fig. S.10(o) presents two instances of this graph family.
@@
    Directed hypercube graphs: Similar to the undirected hypercube graphs of dimension $n$, we consider the vertex set $V(G_{2}^{n})=\{0,1\}^{\times n}$ which contains $N=2^{n}$ vertices. There is a directed edge from vertex $v_{1}$ to $v_{2}$ if the Hamming distance between the vertices is $1$ and the Hamming weight of $v_{1}<v_{2}$. Here, Hamming weight of a vertex refers to the number of $1$s in the corresponding $n$-tuple. This arrangement of directed edges indicates that there is only one source and sink in a directed hypercube graph. The number of edges in a hypercube graph grows as $2^{n-1}n$. Hence, the system size in a directed hypercube graph grows as $N^{\prime}=2^{n}(1+n/2)\sim\mathcal{O}(2^{n}n)$. The number of vertices in the graph assists the graph-family to grow in terms of $2^{n}$, where $n=2,3,\cdots$. Our system size varies $N^{\prime}$ from $8$ to $131072$. Two instances of this graph family can be found in Figure 4(a) of the main manuscript.
S.5.2.2 Random directed graphs
We consider nine random directed graphs in this article. For the sake of reproducibility, we fix seed value of every random directed graph family to be $19$. Further, we followed the constructions from Ref. [28] with slight modifications to avoid bi-directed edges and self-loops in the graphs.
@@
    GN graph: A GN graph grows the network by linking a newly added vertex to one of the existing vertices based on their degree. We define a linear function $f(d)=d$, where $d$ is the degree of the vertex. This function determines the probability of connecting the new vertex to one of the existing vertices. In our work, the value of system size $N^{\prime}$ goes from $9$ to $9997$ for GN graphs with source and sink vertices. Also, the system size varies from $11$ to $9998$ for GN graphs without source and sink vertices. Figures 4(e) and 4(g) of the main manuscript present two instances each for without and with source and sink vertices, respectively.
@@
    GNC graph [38]: A GNC graph is grown by adding a directed edge from a new vertex to an existing vertex chosen randomly. Further, directed edges go from the new vertex to the successors of the existing vertex to which it is connected. A vertex, $v_{u}$, is said to be the successor to a vertex $v_{w}$, if there is a directed edge going from $v_{w}$ to $v_{u}$. For GNC graphs with source and sink vertices, we consider $N^{\prime}$ in the range of $11$ to $8498$. Also, for the family of GNC graphs without any source and sink vertices, the system size $N^{\prime}$ varies from $11$ to $6759$. Two GNC graphs without and with source and sink vertices are depicted in Figs. S.11(a) and S.11(c) respectively.
@@
    GNR graph [37]: The construction of the GNR graphs is similar to the GN graphs. Here, when a new vertex is added, a directed edge appears from the new vertex to a randomly chosen existing vertex, called the target vertex. There is a finite chance for this new directed edge to get redirected to first successor of target vertex. For our work, we set the probability of reconnection to $0.5$. For the GNR graphs with source and sink vertices, the system size grows from $9$ to $9997$. Also, for GNR graphs without source and sink vertices, $N^{\prime}$ varies from $12$ to $10000$. Figs. S.11(e) and S.11(g) respectively present the figures for graphs without and with source and sink vertices belonging to this family.
@@
    Gaussian random partition graphs: The underlying principles for the directed Gaussian random partition graph is similar to its undirected version as described in sub-section S.5.1.2. We follow the construction from Ref. [28] to build directed Gaussian random partition graphs. The values of the parameters to build this graph family are same as the values set in the undirected version. We vary $N^{\prime}$ from $12$ to $10897$ for both the cases. Figure 4(c) contains two instances of this family for without source and sink vertices and Figure S.10(a) has illustrations of two graphs for with source and sink vertices of this family.
@@
    Planted partition graphs: Our construction for the planted partition graph is also similar to their undirected mentioned in the sub-section S.5.1.2. Here, we fix the number of vertices present in each partition to be $5$ and vary the number of partitions to grow the graph. We set the probability of linking vertices inside a partition to $0.8$ and between the partitions to $0.4$. For graphs with sink and source vertices, we vary $N^{\prime}$ from $228$ to $24405$. For graphs without sink and source vertices, we vary $N^{\prime}$ from $228$ to $20208$. Two graphs each for without and with source and sink vertices are presented in Figs. S.10(c) and S.10(e) respectively.
@@
    Navigable small world graphs [34]: The graph is defined on a two-dimensional $n\times n$ grid with $N=n^{2}$ vertices. The distance between two vertices $(a,b)$ and $(c,d)$ is defined as $\mathcal{R}=|a-c|+|b-d|$. Based on the distance between vertices, we define short-range and long-range connections. A vertex $(a,b)$ is connected to all vertices which are at distance $1$ to it. The number of long-range connections of a vertex is set to $1$. The probability of connecting the vertex to a target vertex at a distance $\mathcal{R}$ is proportional to $1/\mathcal{R}^{2}$. For graphs without source and sink vertices, $N^{\prime}$ is varied from $84$ to $20537$, whereas for graphs with source and sink vertices, $N^{\prime}$ goes from $83$ to $14827$. Figs. S.10(g) and S.10(i) show two graphs each for without and with source and sink vertices corresponding to this family.
@@
    Gn,p random graphs: Our construction for the Gn,p random graph is also similar to their counterparts mentioned in the sub-section S.5.1.2 with same values for the parameters. For graphs without sink and source vertices, $N^{\prime}$ is varied from $44$ to $42534$, while for graphs with sink and source vertices, $N^{\prime}$ ranges from $44$ to $21721$. Two graphs each for without and with source and sink vertices can be found in Figs. S.10(k) and S.10(m) respectively.
@@
    Random uniform $k$-out graphs: In this graph, $k$ directed edges go from every vertex of the graph to $k$ vertices chosen randomly without replacement. For our calculations, we set $k=2$. For graphs with source and sink, $N^{\prime}$ takes the values from $12$ to $10943$ and graphs without source and sink, $N^{\prime}$ goes from $13$ to $10817$. Figs. S.10(q) and S.10(s) present two graphs each for without and with source and sink vertices respectively.
@@
    Scale-free graphs [8]: In this graph family, the in-degree and out-degree distributions of vertices follow power laws. We start with a cycle graph with 3 edges. A directed edge from a new vertex to an existing vertex, chosen according to its in-degree distribution, is added with probability $0.41$. An edge between two existing vertices are added with probability $0.54$ and the vertices are chosen according to their in-degree and out-degree distributions. The probability of adding a directed edge from an existing vertex, chosen according to its out-degree distribution, to a new vertex is $0.05$. The probability of choosing an existing vertex would depend on the in-degree (or out-degree) in addition to a constant bias. The constant bias for in-degree distribution is set to be $\delta_{\mathrm{in}}=0.2$ and for out-degree distribution is set to be $\delta_{\mathrm{out}}=0$. Here, we vary the system size $\mathcal{N}$ from $12$ to $10997$ for graphs without sink and source vertices. For graphs with sink and source vertices, $\mathcal{N}$ is varied from $11$ to $9995$. Figs. S.10(u) and S.10(w) present two instances each for graph without and with source and sink vertices respectively.
Appendix S.6 Beyond graph theoretic considerations: Other requirements and challenges
S.6.1  Algorithmic challenges
@@
S.6.2  Quantum hardware challenges
@@
    NISQ era requirements: Although HHL and other QLSs discussed in this work are outside of the scope of NISQ computers and would ideally require a fault-tolerant implementation, it is important to test these solvers on smaller system sizes for various applications even in the NISQ era, continuously pushing the boundaries as quantum hardware advances. A typical theme in carrying out such computations would be to rely upon classical resources (which do not necessarily scale well) to reduce quantum circuit depth and number of qubits, we term these gamut of techniques as resource reduction. In Section VII.1 of the main manuscript, we illustrated these ideas through toy graph instances. Our results show that with current quantum hardware capabilities, we can only probe $(4\times 4)$ size Laplacian matrices (unless we find special graphs and leverage techniques like all-qubit fixing, as explained in Section VII.2 of the main manuscript), indicating the actual gap between what is possible and what is our goal for the long-term. For perspective, it is routine to solve systems of linear equations involving matrices that are about $(10^{9}\times 10^{9})$ in size, for example, in quantum chemical computations on traditional computers [22].
@@
    Fault-tolerant quantum computing era requirements: Of most relevance in the late NISQ/ early fault-tolerant quantum computing eras would be graph families that provide exponential speed-up and polynomial speed-up (with a larger-than-quadratic degree polynomial), in view of overheads associated with quantum error corrected implementations eating up a sub-linear or a lower degree polynomial speed-up offered by good graph families and some better graph families (for example, see Ref. [5], where the authors show that quadratic speedups will not enable quantum advantage on early generation fault-tolerant computers that use surface codes). It is hard to predict the extent to which such issues will plague good and better graph families in the fault-tolerant quantum computing era, where one can use codes with lower thresholds. However, in all of these cases, one aims at reducing either the runtime or the number of physical qubits or in an ideal case, both of them, in fault-tolerant QLS implementations. This would in turn depend on achieving lower physical error rates and faster gate times in the quantum hardware that is used, as well as choice of suitable quantum error correction codes. Each of these is formidable, and until these obstacles are cleared, truly realizing an advantage even with better and best graphs is hard.
Appendix S.7 Some theorems relevant to all-qubit fixing
Theorem S.7.1.
@@
$L\begin{bmatrix}0\\ 0\\ \cdots\\ 0\\ b_{i}\\ 0\\ \cdots\\ 0\\ b_{j}\\ 0\\ \cdots\\ 0\end{bmatrix}=\begin{bmatrix}l_{1i}-l_{1j}\\ l_{2i}-l_{2j}\\ \cdots\\ l_{i-1\ i}-l_{i-1\ j}\\ l_{ii}-l_{ij}\\ l_{i+1\ i}-l_{i+1\ j}\\ \cdots\\ l_{j-1\ i}-l_{j-1\ j}\\ l_{ji}-l_{jj}\\ l_{j+1\ i}-l_{j+1\ j}\\ \cdots\\ l_{Ni}-l_{Nj}\end{bmatrix}$. (S.13)